# DeepBooTS: Dual‑Stream Residual Boosting for Drift‑Resilient Time‑Series Forecasting

### v1

@@ 1 Introduction
Time Series (TS) from natural and engineered systems often evolve under transient conditions (1976_timeseries) and therefore violate stationarity assumptions (hyndman2018forecasting). Classical statistical models such as ARIMA (piccolo1990distance) and exponential smoothing (gardner1985exponential), which depend on fixed statistical properties, struggle to track these shifts (de200625). Deep neural networks have become attractive alternatives because of their powerful nonlinear modeling capabilities (hornik1991approximation), yet even the latest attention‑based (Zhou2021Informer; nie2022time; liu2023itransformer; ye2024frequency; wang2024timemixer; wang2024timexer) and graph‑neural approaches (wu2019graph; yi2024fouriergnn) provide only marginal gains over simple multilayer perceptrons and incur substantial inference overhead (shao2022spatial; zeng2023transformers; yi2024filternet; yu2025drift; yu2025learning).
@@
A deeper issue is that leading architectures like Transformers (NIPS2017_Transformer) and their variants are highly susceptible to concept drift—a mismatch between training and testing distributions. Empirical evidence shows that validation error rises early in training even as training error falls, as shown in Fig. 1, a sign that the learned model fails to generalize as the underlying data distribution shifts. Prevailing solutions, including reversible instance normalization (RevIN) (kim2021reversible) and temporal re‑aggregation, e.g., DLinear (zeng2023transformers) and iTransformer (liu2023itransformer), alleviate mean shifts but leave the variance unstable, which leads to large fluctuations and persistent concept drift. This phenomenon occurs across multiple TS forecasting models and datasets, as shown in Fig. 2(a), underscoring the need for an approach that directly targets the high‑variance component.
@@
We validate DeepBooTS on a wide range of TS benchmarks, including large‑scale TS datasets. As shown in Fig. 2(c), our method consistently outperforms state‑of‑the‑art baselines, achieving an average 15.8% improvement. These results, together with our theoretical analysis, demonstrate that DeepBooTS not only advances the accuracy of TS forecasting but also provides a principled solution to the longstanding problem of concept drift. The code and datasets can be obtained from the appendices we uploaded.
@@ 2.1 Deep Ensemble helps Alleviate Concept Drift
Given a weighted ensemble models $\hat{f}_{\text{ens}}(X)=\frac{1}{L}\sum_{l=1}^{L}\alpha_{l}g_{l}(x)$, where $\sum_{l=1}^{L}\alpha_{l}=1$, we have
$\text{MSE}_{P_{t}}(\hat{f}_{0})-\text{MSE}_{P_{t}}(\hat{f}_{\text{ens}})\approx\frac{(L-1)(1-\alpha_{l})}{L}c^{2}\sigma^{2}_{t}\geq 0.$ (5)
@@ 2.2 Deep Ensemble Model for TS Forecasting
$\ \ \mathbb{E}(Y-\hat{f}_{\text{ens}}(X))^{2}=\mathbb{E}(Y-\sum_{l=1}^{L}\alpha_{l}g_{l}(X))^{2},$ (6)
@@
where $R_{L}$ is the residual term. Further, the aim of the ‘output’ stems is to learn $L$ simple learners in a hierarchy where each learner gives more attention (larger weight) to the hard samples from the previous learner. This is equivalent to the Boosting ensemble learning process, where the final prediction is a weighted sum of $L$ simple learners, and the weights are determined by the previous learner. Let $g_{l}(x)$ denote the $l$-th learner, and $\alpha_{l}$ denote the weight of the $l$-th learner. The overall estimation $\hat{f}_{\text{ens}}(X)$ is a weighted subtraction of the $L$ estimations. For a sample $X$, we have
$\hat{f}_{\text{ens}}(X)=i\sum_{l=0}^{\hbar}\alpha_{2l+1}g_{2l+1}(X)-i\sum_{l=0}^{\hbar}\alpha_{2l}g_{2l}(X),$ (9)
@@
$(Y-\hat{f}_{\text{ens}}(X))^{2}$ $=\sum_{l=0}^{\hbar}i(\alpha_{2l+1}-\alpha_{2l})(Y-g_{l}(X))^{2}$
$-\sum_{l=0}^{\hbar}i(\alpha_{2l+1}-\alpha_{2l})(\hat{f}_{\text{ens}}(X)-g_{l}(X))^{2}$ (10)
@@
The proof is given in Appendix B.5. Clearly, the variance of DeepBooTS is bound by the estimation error (noise error) of each block, and the covariance between blocks. It is evident that the subtraction adopted in DeepBooTS can reduce the variance, thereby mitigating concept drift. On the contrary, switching the aggregation operation of the output stream in the DeepBooTS to addition results in an approximate variance of $\frac{4}{L}\alpha^{2}\nu+3\alpha^{2}\mu$, which is much larger than that of subtraction in Eq. 9. This phenomenon is also supported by our experiments in Section 3.5. Furthermore, Theorem 3 also demonstrates that increasing the number of layers $L$ does not escalate the risk of concept drift. It proves that the deep ensemble models can go deeper, and the test error approaches $\text{Bias}(\hat{f}_{\text{ens}}(X))^{2}+\sigma^{2}$ when $L$ is infinite and doesn’t consider performance-efficient trade-off.
@@ 2.3 Implementation
Equation 12 represents an implicit decomposition of $X$, which differs from the moving average adopted by (wu2021autoformer; zhou2022fedformer; liang2023does) but is similar to (oreshkin2019n). Differently, the residual $R_{l}$ captures what remains unaltered, providing a basis for comparison with the transformed portion.
@@ 3 Experiments
We conduct a comprehensive comparison with the 18 latest SOTA methods, including TimeXer (wang2024timexer), TimeMixer (wang2024timemixer), iTransformer (liu2023itransformer), PatchTST (nie2022time), Crossformer (zhang2022crossformer), SCINet (liu2022scinet), DLinear (zeng2023transformers), FEDformer (zhou2022fedformer) for multivariate forecasting, and FreTS (yi2024frequency), PSLD (liang2024psld), and FourierGNN (yi2024fouriergnn) for large-scale TS datasets, as well as Periodformer (liang2023does), Autoformer (wu2021autoformer), Informer (Zhou2021Informer), LogTrans (li2019enhancing), Reformer (Kitaev2020Reformer), N-BEATS (oreshkin2019n) and N-Hits (challu2023nhits) for univariate forecasting. More experimental settings can be found in Appendix D.
@@ 3.2 Evaluation on Monash TS Datasets
Further, we evaluate the proposed method on 7 Monash TS datasets (godahewa2021monash) (e.g., NN5, M4 and Sunspot, etc.) and 7 diverse metrics (e.g., MAPE, sMAPE, MASE and Quantile, etc.) to systematically evaluate our model. All experiments are compared under the same input length (e.g., II=96) and output lengths (e.g., OO={96, 192, 336 and 720}). As shown in Table 3, the proposed DeepBooTS emerged as the frontrunner, achieving a score of 41 out of 54. Please refer to Appendices Q and R for details about the definition and experimental settings, respectively.
@@ 3.3 Performance on Large-Scale Datasets
The performance comparisons for large-scale TS datasets are summarized in Fig. 5(a), including the CBS dataset with 4,454 nodes (17GB) and the Milano dataset with 10,000 nodes (19GB). On these datasets, it is not feasible to place all nodes on a single GPU for long-term TS forecasting. Therefore, we adopted the random partitioning strategy (liang2024act) for our experiments. Please refer to Appendices C.3 and D for details. Compared to the latest advanced PSLD, the proposed DeepBooTS yields an overall MSE reduction of 8.9% and 6.2% on the CBS and Milano datasets, respectively. In addition, we are pleased to report that our method has been successfully applied to real-world scenarios, e.g., energy and power, showing substantial practical value. For more experiments, please refer to Appendix F.
Figure 6: Ablation studies of DeepBooTS using various Attention. All results are averaged across all prediction lengths. The tick labels of the X-axis are the abbreviation of Attention types. The detailed setup and results are provided in Appendix J.
@@ 4 Conclusion
5 Acknowledgments
This work was supported by Project of Key R&D Program of Shandong Province, China (2025CXGC010107, 2025CXPT095), Taishan Scholars Program: NO.tspd20240814, the Pilot Project for the Integration of Science, Education and Industry of Qilu University of Technology (Shandong Academy of Sciences)(2025ZDZX01), National K&D Development Program of China (2024YFB3312302), and the Qilu Institute of Technology for Talent Project (QIT24TP027), and the Young Talent of Lifting engineering for Science and Technology in Shandong (SDAST2025QTB008).
@@ A.1 Classical Models for TS Forecasting
TS forecasting is a classic research field where numerous methods have been invented to utilize historical series to predict future missing values. Early classical methods (piccolo1990distance; gardner1985exponential) are widely applied because of their well-defined theoretical guarantee and interpretability. For example, ARIMA (piccolo1990distance) initially transforms a non-stationary TS into a stationary one via differencing, and subsequently approximates it using a linear model with several parameters. Exponential smoothing (gardner1985exponential) predicts outcomes at future horizons by computing a weighted average across historical data. In addition, some regression-based methods, e.g., random forest regression (RFR) (liaw2002classification) and support vector regression (SVR) (castro2009online), etc., are also applied to TS forecasting. These methods are straightforward and have fewer parameters to tune, making them a reliable workhorse for TS forecasting. However, their shortcoming is insufficient data fitting ability, especially for high-dimensional series, resulting in limited performance.
@@ A.2 Deep Models for TS Forecasting
The advancement of deep learning has greatly boosted the progress of TS forecasting. Specifically, convolutional neural networks (CNNs) (lecun1998gradient) and recurrent neural networks (RNNs) (connor1994recurrent) have been adopted by many works to model nonlinear dependencies of TS, e.g., LSTNet (lai2018modeling) improve CNNs by adding recursive skip connections to capture long- and short-term temporal patterns; DeepAR (salinas2020deepar) predicts the probability distribution by combining autoregressive methods and RNNs. Several works have improved the series aggregation forms of Attention mechanism, such as operations of exponential intervals adopted in LogTrans (li2019enhancing), ProbSparse activations in Informer (Zhou2021Informer), frequency sampling in FEDformer (zhou2022fedformer) and iterative refinement in Scaleformer (shabani2022scaleformer). Besides, GNNs and Temporal convolutional networks (TCNs) (lea2016temporal) have been utilized in some methods (wu2019graph; li2023dynamic; liu2022scinet; wu2022timesnet) for TS forecasting on graph data. The aforementioned methods solely concentrate on the forms of aggregating input series, overlooking the challenges posed by the concept drift problem.
@@
Since TS exhibit a variety of patterns, it is meaningful and beneficial to decompose them into several components, each representing an underlying category of patterns that evolving over time (1976_timeseries). Several methods, e.g., STL (cleveland1990stl), Prophet (taylor2018forecasting) and N-BEATS (oreshkin2019n), commonly utilize decomposition as a preprocessing phase on historical series. There are also some methods, e.g., Autoformer (wu2021autoformer), FEDformer (zhou2022fedformer), Non-stationary Transformers (liu2022non) and DistPred (liang2025distpred), that harness decomposition into the Attention module. The aforementioned methods attempt to apply decomposition to input series to enhance predictability, reduce computational complexity, or ameliorate the adverse effects of non-stationarity. Nevertheless, these prevalent methods are susceptible to significant concept drift when applied to non-stationary TS.
@@
Furthermore, there are four themes that use deep learning to predict time series: (1) smarter transformers (NIPS2017_Transformer), such as PatchTST (nie2022time), iTransformer (liu2023itransformer) BasisFormer (ni2023basisformer), and TimeXer (wang2024timexer) which restructure attention or add learnable bases to extend context length, cut computation and boost accuracy; (2) competitive non-transformer backbones, including N-HiTS (hierarchical MLP) (challu2023nhits), DLinear (zeng2023transformers), PGN (jia2024pgn) and state-space models like TSMamba (ma2024mamba), TimeMachine (ahamed2024timemachine) and FLDMamba (zhang2025fldmamba), which deliver linear-time inference and rival or surpass transformers on long horizons; (3) foundation-model initiatives, TimeGPT (garza2023timegpt), OneFitAll (zhou2023one), TimeLLM (jin2024timellm), UniTime (liu2024unitime) and DAM (darlowdam2024DAM) that pre-train on massive heterogeneous corpora and achieve impressive zero-shot or few-shot performance across domains; and (4) training and interpretability advances, such as frequency-adaptive normalization (FAN) (ye2024frequency), e.g., FreTS (yi2024frequency), FilterNet (yi2024filternet), and decomposition-aware architectures, e.g., which tackle non-stationarity, quantify uncertainty and make forecasts more transparent.

### v2

@@ 1 Introduction
Time Series (TS) from natural and engineered systems often evolve under transient conditions (Anderson 1976) and therefore violate stationarity assumptions (Hyndman and Athanasopoulos 2018). Classical statistical models such as ARIMA (Piccolo 1990) and exponential smoothing (Gardner Jr 1985), which depend on fixed statistical properties, struggle to track these shifts (De Gooijer and Hyndman 2006). Deep neural networks have become attractive alternatives because of their powerful nonlinear modeling capabilities (Hornik 1991), yet even the latest attention‑based (Zhou et al. 2021; Nie et al. 2022; Liu et al. 2023; Liang et al. 2024b; Liang 2025; Ye et al. 2024; Wang et al. 2024a, b; Liang 2025; Yu et al. 2025a, b) and graph‑neural approaches (Wu et al. 2019; Yi et al. 2024b) provide only marginal gains over simple multilayer perceptrons and incur substantial inference overhead (Shao et al. 2022; Zeng et al. 2023; Yi et al. 2024a).
@@
A deeper issue is that leading architectures like Transformers (Vaswani et al. 2017) and their variants are highly susceptible to concept drift—a mismatch between training and testing distributions. Empirical evidence shows that validation error rises early in training even as training error falls, as shown in Fig. 1, a sign that the learned model fails to generalize as the underlying data distribution shifts. Prevailing solutions, including reversible instance normalization (RevIN) (Kim et al. 2021) and temporal re‑aggregation, e.g., DLinear (Zeng et al. 2023) and iTransformer (Liu et al. 2023), alleviate mean shifts but leave the variance unstable, which leads to large fluctuations and persistent concept drift. This phenomenon occurs across multiple TS forecasting models and datasets, as shown in Fig. 2(a), underscoring the need for an approach that directly targets the high‑variance component.
@@
We validate DeepBooTS on a wide range of TS benchmarks, including large‑scale TS datasets. As shown in Fig. 2(c), our method consistently outperforms state‑of‑the‑art baselines, achieving an average 15.8% improvement. These results, together with our theoretical analysis, demonstrate that DeepBooTS not only advances the accuracy of TS forecasting but also provides a principled solution to the longstanding problem of concept drift.
@@ 2.1 Deep Ensemble helps Alleviate Concept Drift
Given a weighted ensemble models $\hat{f}(X)=\frac{1}{L}\sum_{l=1}^{L}\alpha_{l}g_{l}(x)$, where $\sum_{l=1}^{L}\alpha_{l}=1$, we have
$\text{MSE}_{P_{t}}(\hat{f}_{0})-\text{MSE}_{P_{t}}(\hat{f})\approx\frac{(L-1)(1-\alpha_{l})}{L}c^{2}\sigma^{2}_{t}\geq 0.$ (5)
@@ 2.2 Deep Ensemble Model for TS Forecasting
$\ \ \mathbb{E}(Y-\hat{f}(X))^{2}=\mathbb{E}(Y-\sum_{l=1}^{L}\alpha_{l}g_{l}(X))^{2},$ (6)
@@
where $R_{L}$ is the residual term. Further, the aim of the ‘output’ stems is to learn $L$ simple learners in a hierarchy where each learner gives more attention (larger weight) to the hard samples from the previous learner. This is equivalent to the Boosting ensemble learning process, where the final prediction is a weighted sum of $L$ simple learners, and the weights are determined by the previous learner. Let $g_{l}(x)$ denote the $l$-th learner, and $\alpha_{l}$ denote the weight of the $l$-th learner. The overall estimation $\hat{f}(X)$ is a weighted subtraction of the $L$ estimations. For a sample $X$, we have
$\hat{f}(X)=i\sum_{l=0}^{\hbar}\alpha_{2l+1}g_{2l+1}(X)-i\sum_{l=0}^{\hbar}\alpha_{2l}g_{2l}(X),$ (9)
@@
$(Y-\hat{f}(X))^{2}$ $=\sum_{l=0}^{\hbar}i(\alpha_{2l+1}-\alpha_{2l})(Y-g_{l}(X))^{2}$
$-\sum_{l=0}^{\hbar}i(\alpha_{2l+1}-\alpha_{2l})(\hat{f}(X)-g_{l}(X))^{2}$ (10)
@@
The proof is given in Appendix B.5. Clearly, the variance of DeepBooTS is bound by the estimation error (noise error) of each block, and the covariance between blocks. It is evident that the subtraction adopted in DeepBooTS can reduce the variance, thereby mitigating concept drift. On the contrary, switching the aggregation operation of the output stream in the DeepBooTS to addition results in an approximate variance of $\frac{4}{L}\alpha^{2}\nu+3\alpha^{2}\mu$, which is much larger than that of subtraction in Eq. 9. This phenomenon is also supported by our experiments in Section 3.5. Furthermore, Theorem 3 also demonstrates that increasing the number of layers $L$ does not escalate the risk of concept drift. It proves that the deep ensemble models can go deeper, and the test error approaches $\text{Bias}(\hat{f}(X))^{2}+\sigma^{2}$ when $L$ is infinite and doesn’t consider performance-efficient trade-off.
@@ 2.3 Implementation
Equation 12 represents an implicit decomposition of $X$, which differs from the moving average adopted by (Wu et al. 2021; Zhou et al. 2022; Liang et al. 2023) but is similar to (Oreshkin et al. 2019). Differently, the residual $R_{l}$ captures what remains unaltered, providing a basis for comparison with the transformed portion.
@@ 3 Experiments
We conduct a comprehensive comparison with the 18 latest SOTA methods, including TimeXer (Wang et al. 2024b), TimeMixer (Wang et al. 2024a), iTransformer (Liu et al. 2023), PatchTST (Nie et al. 2022), Crossformer (Zhang and Yan 2022), SCINet (Liu et al. 2022a), DLinear (Zeng et al. 2023), FEDformer (Zhou et al. 2022) for multivariate forecasting, and FreTS (Yi et al. 2024c), PSLD (Liang et al. 2024d), and FourierGNN (Yi et al. 2024b) for large-scale TS datasets, as well as Periodformer (Liang et al. 2024c), Autoformer (Wu et al. 2021), Informer (Zhou et al. 2021), LogTrans (Li et al. 2019), Reformer (Kitaev et al. 2020), N-BEATS (Oreshkin et al. 2019) and N-Hits (Challu et al. 2023) for univariate forecasting. More experimental settings can be found in Appendix D.
@@ 3.2 Evaluation on Monash TS Datasets
Further, we evaluate the proposed method on 7 Monash TS datasets (Godahewa et al. 2021) (e.g., NN5, M4 and Sunspot, etc.) and 7 diverse metrics (e.g., MAPE, sMAPE, MASE and Quantile, etc.) to systematically evaluate our model. All experiments are compared under the same input length (e.g., II=96) and output lengths (e.g., OO={96, 192, 336 and 720}). As shown in Table 3, the proposed DeepBooTS emerged as the frontrunner, achieving a score of 41 out of 54. Please refer to Appendices Q and R for details about the definition and experimental settings, respectively.
@@ 3.3 Performance on Large-Scale Datasets
The performance comparisons for large-scale TS datasets are summarized in Fig. 5(a), including the CBS dataset with 4,454 nodes (17GB) and the Milano dataset with 10,000 nodes (19GB). On these datasets, it is not feasible to place all nodes on a single GPU for long-term TS forecasting. Therefore, we adopted the random partitioning strategy (Liang et al. 2024a) for our experiments. Please refer to Appendices C.3. and D for details. Compared to the latest advanced PSLD, the proposed DeepBooTS yields an overall MSE reduction of 8.9% and 6.2% on the CBS and Milano datasets, respectively. In addition, we are pleased to report that our method has been successfully applied to real-world scenarios, e.g., energy and power, showing substantial practical value. For more experiments, please refer to Appendix F.
Figure 6: Ablation studies of DeepBooTS using various Attention. All results are averaged across all prediction lengths. The tick labels of the X-axis are the abbreviation of Attention types. The detailed setup and results are provided in Appendix Q.
@@ 4 Conclusion
Acknowledgments
This work was supported by Project of Key R&D Program of Shandong Province, China (2025CXGC010107, 2025CXPT095), Taishan Scholars Program: NO.tspd20240814, the Pilot Project for the Integration of Science, Education and Industry of Qilu University of Technology (Shandong Academy of Sciences)(2025ZDZX01), National Key R&D Program of China (2024YFB3312302), and the Qilu Institute of Technology for Talent Project (QIT24TP027), and the Young Talent of Lifting engineering for Science and Technology in Shandong (SDAST2025QTB008).
@@ A.1 Classical Models for TS Forecasting
TS forecasting is a classic research field where numerous methods have been invented to utilize historical series to predict future missing values. Early classical methods (Piccolo 1990; Gardner Jr 1985) are widely applied because of their well-defined theoretical guarantee and interpretability. For example, ARIMA (Piccolo 1990) initially transforms a non-stationary TS into a stationary one via differencing, and subsequently approximates it using a linear model with several parameters. Exponential smoothing (Gardner Jr 1985) predicts outcomes at future horizons by computing a weighted average across historical data. In addition, some regression-based methods, e.g., random forest regression (RFR) (Liaw et al. 2002) and support vector regression (SVR) (Castro-Neto et al. 2009), etc., are also applied to TS forecasting. These methods are straightforward and have fewer parameters to tune, making them a reliable workhorse for TS forecasting. However, their shortcoming is insufficient data fitting ability, especially for high-dimensional series, resulting in limited performance.
@@ A.2 Deep Models for TS Forecasting
The advancement of deep learning has greatly boosted the progress of TS forecasting. Specifically, convolutional neural networks (CNNs) (LeCun et al. 1998) and recurrent neural networks (RNNs) (Connor et al. 1994) have been adopted by many works to model nonlinear dependencies of TS, e.g., LSTNet (Lai et al. 2018) improve CNNs by adding recursive skip connections to capture long- and short-term temporal patterns; DeepAR (Salinas et al. 2020) predicts the probability distribution by combining autoregressive methods and RNNs. Several works have improved the series aggregation forms of Attention mechanism, such as operations of exponential intervals adopted in LogTrans (Li et al. 2019), ProbSparse activations in Informer (Zhou et al. 2021), frequency sampling in FEDformer (Zhou et al. 2022) and iterative refinement in Scaleformer (Shabani et al. 2022). Besides, GNNs and Temporal convolutional networks (TCNs) (Lea et al. 2016) have been utilized in some methods (Wu et al. 2019; Li et al. 2023; Liu et al. 2022a; Wu et al. 2022a) for TS forecasting on graph data. The aforementioned methods solely concentrate on the forms of aggregating input series, overlooking the challenges posed by the concept drift problem.
@@
Since TS exhibit a variety of patterns, it is meaningful and beneficial to decompose them into several components, each representing an underlying category of patterns that evolving over time (Anderson 1976). Several methods, e.g., STL (Cleveland et al. 1990), Prophet (Taylor and Letham 2018) and N-BEATS (Oreshkin et al. 2019), commonly utilize decomposition as a preprocessing phase on historical series. There are also some methods, e.g., Autoformer (Wu et al. 2021), FEDformer (Zhou et al. 2022), Non-stationary Transformers (Liu et al. 2022b), Minusformer (Liang et al. 2024b) and DistPred (Liang 2025), that harness decomposition into the Attention module. The aforementioned methods attempt to apply decomposition to input series to enhance predictability, reduce computational complexity, or ameliorate the adverse effects of non-stationarity. Nevertheless, these prevalent methods are susceptible to significant concept drift when applied to non-stationary TS.
@@
Furthermore, there are four themes that use deep learning to predict time series: (1) smarter transformers (Vaswani et al. 2017), such as PatchTST (Nie et al. 2022), iTransformer (Liu et al. 2023) BasisFormer (Ni et al. 2023), and TimeXer (Wang et al. 2024b) which restructure attention or add learnable bases to extend context length, cut computation and boost accuracy; (2) competitive non-transformer backbones, including N-HiTS (hierarchical MLP) (Challu et al. 2023), DLinear (Zeng et al. 2023), PGN (Jia et al. 2024) and state-space models like TSMamba (Ma et al. 2024), TimeMachine (Ahamed and Cheng 2024) and FLDMamba (Zhang et al. 2025), which deliver linear-time inference and rival or surpass transformers on long horizons; (3) foundation-model initiatives, TimeGPT (Garza et al. 2023), OneFitAll (Zhou et al. 2023), TimeLLM (Jin et al. 2024), UniTime (Liu et al. 2024) and DAM (Darlow et al. 2024) that pre-train on massive heterogeneous corpora and achieve impressive zero-shot or few-shot performance across domains; and (4) training and interpretability advances, such as frequency-adaptive normalization (FAN) (Ye et al. 2024), e.g., FreTS (Yi et al. 2024c), FilterNet (Yi et al. 2024a), and decomposition-aware architectures, e.g., which tackle non-stationarity, quantify uncertainty and make forecasts more transparent.